# FRNet: Frustum-Range Networks for Scalable LiDAR Segmentation

### v2

@@ Abstract
LiDAR segmentation has become a crucial component in advanced autonomous driving systems. Recent range-view LiDAR segmentation approaches show promise for real-time processing. However, they inevitably suffer from corrupted contextual information and rely heavily on post-processing techniques for prediction refinement. In this work, we propose FRNet, a simple yet powerful method aimed at restoring the contextual information of range image pixels using corresponding frustum LiDAR points. Firstly, a frustum feature encoder module is used to extract per-point features within the frustum region, which preserves scene consistency and is crucial for point-level predictions. Next, a frustum-point fusion module is introduced to update per-point features hierarchically, enabling each point to extract more surrounding information via the frustum features. Finally, a head fusion module is used to fuse features at different levels for final semantic prediction. Extensive experiments conducted on four popular LiDAR segmentation benchmarks under various task setups demonstrate the superiority of FRNet. Notably, FRNet achieves 73.3% and 82.5% mIoU scores on the testing sets of SemanticKITTI and nuScenes. While achieving competitive performance, FRNet operates 5 times faster than state-of-the-art approaches. Such high efficiency opens up new possibilities for more scalable LiDAR segmentation. The code has been made publicly available at https://github.com/Xiangxu-0103/FRNet.
@@ I Introduction
LiDAR segmentation, an crucial and indispensable component in modern autonomous driving, robotics, and other safety-critical applications, has witnessed substantial advancements recently [1, 2, 3, 4]. The principal challenge lies in devising a scalable LiDAR segmentation system that strikes a delicate balance between efficiency and accuracy, especially in resource-constrained operational scenarios [5, 6].
Existing LiDAR segmentation methods adopt various data representation perspectives, each with its own set of trade-offs in terms of accuracy and computational efficiency. The point-based methods [7, 8, 9, 10, 11, 12] manipulate the original point cloud to preserve raw spatial information, but entail computationally expensive neighborhood searches to construct local structures. This demand results in considerable computation overhead and limits their capability to deal with large-scale scene point clouds [13, 14, 10]. Sparse-voxel-based methods [15, 16, 17, 18] transform scattered LiDAR points into regular voxel representations and leverage popular sparse convolutions to extract features. However, this process also demands heavy computations, especially for applications that require high voxel resolutions [16, 15, 19]. Multi-view methods [20, 21, 22, 23] extract features from various representations to enhance prediction accuracy [24, 25, 26]. While this strategy can improve performance, it aggregates the computation required for all representations. Albeit these representation methods achieved satisfactory performance, their computation overhead hinders them from real-time online applications in real life [27, 23].
@@
Observing the above issues, in this work, we propose a scalable Frustum-Range Network (FRNet) for LiDAR segmentation, which incorporates points into the range image, achieving a superior balance between efficiency and accuracy. FRNet consists of three main components. Firstly, a frustum feature encoder (FFE) is utilized to group all points with the same frustum region into corresponding range-view pixels in relation to the range image using multiple multi-layer perceptrons (MLPs). This allows for the preservation of all points and the prediction of semantic labels for these points in an end-to-end manner. Subsequently, the point features are pooled to represent the frustum region and formatted into a 2D representation, which is then subjected to traditional convolutions. Secondly, a frustum-point (FP) fusion module is employed to efficiently update the hierarchical features of each point during each convolutional stage. This module includes frustum-to-point fusion to update per-point features and point-to-frustum fusion to enhance frustum features. As a result, all points extract larger local features based on the frustum region. Finally, a fusion head (FH) module is designed to leverage features from different levels to generate individual features for each point, facilitating end-to-end prediction without the need for post-processing techniques. As shown in Fig. 1, the proposed FRNet achieves great improvement among range-view methods while still maintaining high efficiency.
Data augmentation plays an important role in training robust and generalizable models. Inspired by the success of mixing strategies [44, 45, 46], we introduce a novel mixing method dubbed FrustumMix, which operates based on frustum units. Specifically, we randomly divide point cloud scenes into several frustum regions along inclination and azimuth directions and swap the corresponding regions from different scenes, with the neighborhood of each frustum region coming from the other scene. FrustumMix provides a straightforward and lightweight method to adapt FRNet by focusing on the frustum region. Furthermore, to address the issue of empty pixels in the 2D representation, we propose RangeInterpolation to reconstruct the semantic surface based on surrounding point information. It first transforms the LiDAR point cloud into a range image, as done in previous works [29, 30, 31, 32]. For empty pixels in the range image, we then aggregate surrounding range information within a pre-defined window using an average pooling operation to generate a new point. RangeInterpolation is able to assist in creating a more compact 2D representation with fewer empty pixels, enabling convolutions to learn more semantically coherent information.
In summary, the key contributions of this work can be listed as follows:
@@
  • We propose two novel 3D data augmentation techniques to assist in training robust and generalizable models. FrustumMix generates more complex scenes by mixing two different scans, while RangeInterpolation reconstructs the semantic surfaces based on surrounding range information. These techniques help generate a compact 2D representation and learn more context-aware features.
@@ II-A Scalable LiDAR Segmentation
Although point-view [7, 8, 10, 11, 12], sparse-voxel-view [15, 16, 17, 18], and multi-view [20, 21, 22, 23, 47, 48] methods have achieved significant success in LiDAR segmentation, they often suffer from high computational costs and are unable to achieve real-time processing. Recently, various methods have been proposed to project LiDAR points into range images along inclination and azimuth directions to balance the efficiency and accuracy of segmentation [49]. SqueezeSeg [29] and SqueezeSegV2 [50] use the lightweight model SqueezeNet [51] to retain information. SqueezeSegV3 [52] introduces a Spatially-Adaptive Convolution module to adjust kernel weights based on the locations of range images. RangeNet++ [30] integrates DarkNet into SqueezeSeg [29] and proposes an efficient KNN post-processing for segmentation. RangeViT [31] and RangeFormer [32] introduce transformer blocks to extract both local and global information from the range image. However, these methods only involve projected points and fail to extract contextual information for discarded points. To address this problem, WaffleIron [11] proposes to mix point and image features with residual connections. Although achieving promising performance, it requires downsampling points based on voxel grid and space neighborhood search for point feature embedding, which hinders it from real-time processing and end-to-end predicting over the entire point cloud. In this work, we propose to integrate point and image features efficiently, preserving both 3D geometric information from raw point features and efficiency from 2D convolutions, without the need for any pre- or post-processing techniques, which yields superior scalability for real-time LiDAR segmentation.
@@ II-B LiDAR Data Augmentation
Recent works have explored various data augmentation techniques on point clouds [53, 54], which have shown promise in improving performance for indoor scene understanding but exhibit limited generalization to outdoor scenes. Mix3D [46] creates new training samples by combining two scenes out of context, which requires a larger memory cost. PolarMix [45] proposes to mix two LiDAR scenes based on azimuth angles and cropping instances from one scene to another, while LaserMix [44] suggests exchanging labeled scans with unlabeled scans along the inclination direction for efficient semi-supervised learning. RangeFormer [32] presents a lightweight augmentation method that operates only on projected points in range images, disregarding occluded points. In this work, we introduce augmentation techniques aimed at training a robust and scalable model for LiDAR segmentation, including FrustumMix and RangeInterpolation.
@@ II-C LiDAR Post-Processing
Prior works have proposed various post-processing techniques to reconstruct the semantic labels of occluded points based on 2D results. SqueezeSeg [29] employs the traditional conditional random field (CRF) as a recurrent neural network (RNN). RangeNet++ [30] introduces a fast K-Nearest-Neighbor (KNN) search to aggregate final results across the entire point cloud. FIDNet [55] uses a Nearest Label Assignment (NLA) method for label smoothing. RangeFormer [32] divides full point clouds into sub-clouds and infers each subset in a supervised manner. However, these methods heavily rely on 2D results and fail to extract point features over entire points in a learnable manner. In contrast, KPRNet [56] and RangeViT [31] incorporate a KPConv [7] block to extract point features in an end-to-end manner. Nevertheless, this approach requires substantial computational resources to search neighborhoods in 3D space, which hinders them from real-time processing. In this work, we partition points into frustum regions and extract local features of points in a 2D manner, avoiding time-consuming neighbor searches. Based on the learned point features, FRNet predicts results over the entire point cloud in an end-to-end manner with high efficiency.
@@ III-A Pilot Study
LiDAR segmentation aims to assign semantic labels to each point in the point cloud [59]. A common approach is to operate directly on the LiDAR points without any pre-processing, following the paradigm of PointNet++ [60]. However, point-based operations often require extensive computations and can yield sub-optimal performance due to the sparsity of outdoor scenes. To address this problem, many recent works have proposed projecting LiDAR points onto a regular intermediate representation, known as range images, based on spherical coordinates to enhance efficiency. Specifically, given a point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}$ with $N$ points, where $\mathbf{p}_{i}\in\mathbb{R}^{3+L}$ includes the coordinate $(x_{i},y_{i},z_{i})$ and additional $L$-dimensional features (intensity, elongation, etc.), we map the point to its corresponding 2D position $(u_{i},v_{i})$ in the range image via the following transformation:
$\binom{u_{i}}{v_{i}}=\binom{\frac{1}{2}[1-\arctan(y_{i},x_{i})\pi^{-1}]W}{[1-(\arcsin(z_{i}d_{i}^{-1})+\phi_{\text{down}})f^{-1}H]}~{},$ (1)
where $d_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}+z_{i}^{2}}$ is the depth of the point; $\phi=|\phi_{\text{up}}|+|\phi_{\text{down}}|$ represents the vertical field-of-view (FOV) of the sensor, and $\phi_{\text{up}}$ are $\phi_{\text{down}}$ are the inclination angles at the upward and downward directions, respectively; $H$ and $W$ are the height and width of the range image. However, such projection often results in the loss of occluded points and requires post-processing techniques to reconstruct the entire semantic information [32].
We conducted a pilot study on various range-view methods, including RangeNet++ [30], SalsaNext [57], FIDNet [55], and CENet [58]. We compared their performances on the original points (i.e., after the 2D-to-3D projection) using the widely-used KNN post-processing with various nearest neighbor parameters, as well as their performance on the 2D range images. As illustrated in Fig. 2, the post-processing procedures are often unsupervised, and the choice of hyperparameters can significantly impact the final performance. To address this issue, we propose a scalable Frustum-Range Network (FRNet) capable of directly predicting semantic labels for each point while maintaining high efficiency.
@@ III-B Frustum-Range Architecture
Problem Definition. Given a point cloud $\mathcal{P}$ acquired by the LiDAR sensor, the objective of FRNet is to employ a feed-forward network $\mathcal{G}$ to predict semantic labels $\mathcal{\hat{Y}}$ for each individual point as follows:
$\mathcal{\hat{Y}}=\mathcal{G}(\mathcal{P},\theta)~{},$ (2)
@@
Frustum Feature Encoder. In LiDAR segmentation, the accurate prediction of semantic labels for each point requires the extraction of individual features for each point. Preceding the conversion of the point cloud into a 2D representation, the frustum feature encoder plays a pivotal role in extracting per-point features through a series of MLPs. This process is crucial for accurate label prediction. Firstly, We group the point cloud to yield a set of $M$ frustum regions $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{M}\}$ according to Eq. 1, where $\mathcal{P}_{i}\in\mathbb{R}^{N_{i}\times(3+L)}$, consisting of $N_{i}$ points. To explicitly embed the frustum structure, we leverage an average pooling function to obtain the cluster points $\widetilde{\mathcal{P}}=\{\widetilde{\mathcal{P}}_{1},\widetilde{\mathcal{P}}_{2},...,\widetilde{\mathcal{P}}_{M}\}$, where $\widetilde{\mathcal{P}}_{i}\in\mathbb{R}^{1\times(3+L)}$. Subsequently, we construct a frustum graph between the point cloud $\mathcal{P}$ and the cluster points $\widetilde{\mathcal{P}}$ within each frustum region. The per-point features are obtained as follows:
@@
Frustum-Point Fusion Module. Following prior works [55, 58, 32], we utilize a 2D backbone composed of multiple convolution blocks that efficiently extract hierarchical frustum features. Each stage block is accompanied by a frustum-point fusion module, facilitating the incremental update of point and frustum features. As illustrated in Fig. 4, the frustum-point fusion module consists of two essential components: 1) A frustum-to-point fusion that enables the hierarchical update of per-point features, and 2) A point-to-frustum fusion that fuses individual features of each point within its corresponding frustum region. The $i$-th frustum-point fusion module takes $\mathcal{F}_{p}^{i-1}$ and ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ as inputs, where ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is the output of the 2D convolution block that takes $\mathcal{F}_{f}^{i-1}$ as input. To efficiently extract context information of the point, we employ a frustum-to-point fusion block to integrate local features of the frustum region into points. Firstly, ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is back-projected to the point-level features according to the projection index. Subsequently, the back-projected features are concatenated with $\mathcal{F}_{p}^{i-1}$ and passed through an MLP to update per-point features, avoiding the high-computation in neighborhood searches:
$\mathcal{F}_{p}^{i}=\text{MLP}([\text{Inflat}(\widetilde{\mathcal{F}}_{f}^{i-1});\mathcal{F}_{p}^{i-1}])~{}.$ (4)
As discussed above, the 2D representation is non-compact due to the sparsity of the point cloud. The convolution block inevitably introduces noisy information for the empty frustum, leading to compact feature representation. To preserve the sparsity attributes of the 2D representation, we design a point-to-frustum fusion module to fuse the sparse and dense frustum features. Specifically, the updated point features $\mathcal{F}_{p}^{i}$ are first pooled into the corresponding frustum region to yield a non-compact 2D representation. Subsequently, the non-compact frustum feature is concatenated with the compact frustum feature ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ and passed through a simple convolutional layer to reduce the number of channels:
$\mathcal{F}_{\text{fuse}}^{i}=\text{Conv}([\text{Flat}(\mathcal{F}_{p}^{i});\widetilde{\mathcal{F}}_{f}^{i-1}])~{},$ (5)
where $\text{Conv}(\cdot)$ represents a convolution layer with batch normalization and activation function. Finally, the fused features are utilized to further enhance the frustum features in a residual-attentive manner:
$\mathcal{F}_{f}^{i}=\widetilde{\mathcal{F}}_{f}^{i-1}+\sigma(h(\mathcal{F}_{\text{fuse}}^{i},\theta))\odot\mathcal{F}_{\text{fuse}}^{i}~{},$ (6)
@@
Fusion Head Module. FRNet aims to assign a category for each point in the point cloud, which requires point-level features for end-to-end label prediction. Although output point features at the last layer of the frustum-point fusion module can yield promising performance, we observe this is not the optimal solution. Features from different levels have different context-aware information. Lower-level features are geometric while higher-level features are semantically rich. Leveraging these different level features can enhance performance. Drawing inspiration from FIDNet [55] and CENet [58], we first combine all point and frustum features from each stage of the backbone and fuse them to reduce channel dimension:
@@
where $V$ is the number of stages in the backbone. Fusing backbone features with different receptive fields helps obtain more context-aware and semantic information. Additionally, point features $\mathcal{F}_{p}^{\text{out}}$ contains local information from neighborhood frustum regions, while frustum features $\mathcal{F}_{f}^{\text{out}}$ contains global information within the corresponding frustum region. Meanwhile, $\mathcal{F}_{p}^{0}$ from the frustum feature encoder module constructs graph structures within the frustum region and between neighborhood frustum regions. Progressively fusing these features helps extract features from local to global, and from geometric to semantic, resulting in more accurate predictions. To fuse $\mathcal{F}_{f}^{\text{out}}$ and $\mathcal{F}_{p}^{\text{out}}$, we first back-project frustum features to point-level and use an MLP layer ensure the same channels as $\mathcal{F}_{p}^{\text{out}}$. Then, an addition operation is applied to fuse the two features. The same operation is utilized to fuse the new features with $\mathcal{F}_{p}^{0}$. The overall operation can be formulated as:
@@
Optimization. In the conventional setup, a point-level loss function is commonly employed to supervise the final semantic scores. However, in this work, we propose optimizing frustum features with additional auxiliary heads, utilizing the frustum-level loss function. This approach requires generating pseudo frustum labels. Typically, a frustum region encompasses multiple points, which may not all belong to the same category. To address this, we adopt a statistical approach inspired by sparse-voxel methods [16]. We generate a pseudo label for each frustum by first counting the frequency of each category existing in the given frustum region, and then choosing the category with the highest frequency as the pseudo label of the frustum. Finally, the frustum-level features are supervised using the generated pseudo labels. The overall loss function can be formulated as follows:
@@
where $\mathcal{L}_{p}$ and $\mathcal{L}_{f}$ represent the point-level and frustum-level loss functions, respectively, and $\lambda$ is a hyperparameter that controls the weight of frustum supervision. $\mathcal{L}_{p}$ is calculated by cross-entropy loss, while $\mathcal{L}_{f}$ consists of cross-entropy loss, Lovasz-Softmax loss [61], and boundary loss [62].

### v3

@@ Abstract
LiDAR segmentation has become a crucial component of advanced autonomous driving systems. Recent range-view LiDAR segmentation approaches show promise for real-time processing. However, they inevitably suffer from corrupted contextual information and rely heavily on post-processing techniques for prediction refinement. In this work, we propose FRNet, a simple yet powerful method aimed at restoring the contextual information of range image pixels using corresponding frustum LiDAR points. First, a frustum feature encoder module is used to extract per-point features within the frustum region, which preserves scene consistency and is critical for point-level predictions. Next, a frustum-point fusion module is introduced to update per-point features hierarchically, enabling each point to extract more surrounding information through the frustum features. Finally, a head fusion module is used to fuse features at different levels for final semantic predictions. Extensive experiments conducted on four popular LiDAR segmentation benchmarks under various task setups demonstrate the superiority of FRNet. Notably, FRNet achieves 73.3% and 82.5% mIoU scores on the testing sets of SemanticKITTI and nuScenes. While achieving competitive performance, FRNet operates 5 times faster than state-of-the-art approaches. Such high efficiency opens up new possibilities for more scalable LiDAR segmentation. The code has been made publicly available at https://github.com/Xiangxu-0103/FRNet.
@@ I Introduction
LiDAR segmentation, a crucial and indispensable component in modern autonomous driving, robotics, and other safety-critical applications, has witnessed substantial advancements recently [1, 2, 3, 4]. The principal challenge lies in devising a scalable LiDAR segmentation system that strikes a delicate balance between efficiency and accuracy, especially in resource-constrained operational scenarios [5, 6]. Here, “scalable” refers to the system’s ability to maintain high performance while managing increased data loads without compromising real-time processing capabilities.
Existing LiDAR segmentation methods adopt various data representation perspectives, each with inherent trade-offs in terms of accuracy and computational efficiency. Point-based methods [7, 8, 9, 10, 11, 12] manipulate the original point cloud to preserve spatial granularity but entail computationally expensive neighborhood searches to construct local structures. This demand results in significant computational overhead and limits their capability to deal with large-scale point clouds [13, 14, 10]. Sparse-voxel-based methods [15, 16, 17, 18, 19] transform scattered LiDAR points into regular voxel grids and leverage popular sparse convolutions to extract features. However, this process also demands heavy computations, especially for applications that require high voxel resolutions [16, 15, 20]. Multi-view methods [21, 22, 23, 24] extract features from various representations to enhance prediction accuracy [25, 26, 27]. While this strategy can improve performance, it aggregates the computation required for all representations. Albeit these representation methods achieved satisfactory performance, their computation overhead hinders them from real-time online applications in real life [28, 24, 29, 30, 31].
@@
Observing the above issues, in this work, we propose a Frustum-Range Network (FRNet) for scalable LiDAR segmentation, which incorporates points into the range image, achieving a superior balance between efficiency and accuracy. FRNet consists of three main components. Firstly, a Frustum Feature Encoder (FFE) is utilized to group all points with the same frustum region into corresponding range-view pixels in relation to the range image using multiple multi-layer perceptrons (MLPs). This allows for the preservation of all points and the prediction of semantic labels for these points in an end-to-end manner. Subsequently, the point features are pooled to represent the frustum region and formatted into a 2D representation, which is then subjected to traditional convolutions. Secondly, a Frustum-Point (FP) fusion module is employed to efficiently update the hierarchical features of each point during each convolutional stage. This module includes frustum-to-point fusion to update per-point features and point-to-frustum fusion to enhance frustum features. As a result, all points extract larger local features based on the frustum region. Finally, a Fusion Head (FH) module is designed to leverage features from different levels to generate individual features for each point, facilitating end-to-end prediction without the need for post-processing techniques. As shown in Fig. 1, the proposed FRNet achieves great improvement among range-view methods while still maintaining high efficiency.
Data augmentation plays an important role in training robust and generalizable models. Inspired by the success of mixing strategies [48, 49, 50], we introduce a fine-grained mixing method dubbed FrustumMix, which operates based on frustum units. Specifically, we randomly divide point cloud scenes into several frustum regions along inclination and azimuth directions and swap the corresponding regions from different scenes, with the neighborhood of each frustum region coming from the other scene. FrustumMix provides a straightforward and lightweight method to adapt FRNet by focusing on the frustum region. Furthermore, to address the issue of empty pixels in the 2D representation, we propose RangeInterpolation to reconstruct the semantic surface based on surrounding point information. It first transforms the LiDAR point cloud into a range image, as done in previous works [33, 34, 35, 36]. For empty pixels in the range image, we then aggregate surrounding range information within a pre-defined window using an average pooling operation to generate a new point. RangeInterpolation is able to assist in creating a more compact 2D representation with fewer empty pixels, enabling convolutions to learn more semantically coherent information.
The key contributions of this work are summarized as:
@@
  • We propose two 3D data augmentation techniques to assist in training robust and generalizable models. FrustumMix generates more complex scenes by mixing two different scans, while RangeInterpolation reconstructs the semantic surfaces based on surrounding range information. These techniques help generate a compact 2D representation and learn more context-aware features.
@@ II-A Scalable LiDAR Segmentation
Although point-view [7, 8, 10, 11, 12], sparse-voxel-view [15, 16, 17, 18], and multi-view [21, 22, 23, 24, 51, 52] methods have achieved significant success in LiDAR segmentation, they often suffer from high computational costs and are unable to achieve real-time processing. Recently, various methods have been proposed to project LiDAR points into range images along inclination and azimuth directions to balance the efficiency and accuracy of segmentation [53]. SqueezeSeg [33] and SqueezeSegV2 [54] use the lightweight model SqueezeNet [55] to retain information. SqueezeSegV3 [56] introduces a Spatially-Adaptive Convolution module to adjust kernel weights based on the locations of range images. RangeNet++ [34] integrates DarkNet into SqueezeSeg [33] and proposes an efficient KNN post-processing for segmentation. RangeViT [35] and RangeFormer [36] introduce transformer blocks to extract both local and global information from the range image. However, these methods only involve projected points and fail to extract contextual information for discarded points. To address this problem, WaffleIron [11] proposes to mix point and image features with residual connections. Although achieving promising performance, it requires downsampling points based on voxel grid and space neighborhood search for point feature embedding, which hinders it from real-time processing and end-to-end predictions over the entire point cloud. In this work, we propose an efficient integration of point and image features that preserves both the 3D geometric information from raw point features and the efficiency of 2D convolutions, eliminating the need for pre- or post-processing techniques and yielding superior scalability for real-time LiDAR segmentation.
@@ II-B LiDAR Data Augmentation
Recent works have explored various data augmentation techniques for point clouds [57, 58], which have shown promise in improving performance for indoor scene understanding but exhibit limited generalization to outdoor scenes. Mix3D [50] creates new training samples by combining two out-of-context scenes, which incurs higher memory costs. PolarMix [49] proposes to mix two LiDAR scenes based on azimuth angles and crop instances from one scene to another, while LaserMix [48] suggests exchanging labeled scans with unlabeled scans along the inclination direction for efficient semi-supervised learning. RangeFormer [36] presents a lightweight augmentation method that operates only on projected points in range images, disregarding occluded points. In this work, we introduce augmentation techniques aimed at training a robust and scalable model for LiDAR segmentation, including FrustumMix and RangeInterpolation.
@@ II-C LiDAR Post-Processing
Prior works have proposed various post-processing techniques to reconstruct the semantic labels of occluded points based on 2D results. SqueezeSeg [33] employs the traditional conditional random field (CRF) as a recurrent neural network (RNN). RangeNet++ [34] introduces a fast K-Nearest-Neighbor (KNN) search to aggregate final results across the entire point cloud. FIDNet [59] uses a Nearest Label Assignment (NLA) method for label smoothing. RangeFormer [36] divides full point clouds into sub-clouds and infers each subset in a supervised manner. However, these methods heavily rely on 2D results and fail to extract point features over the entire point cloud in a supervised manner. In contrast, KPRNet [60] and RangeViT [35] incorporate a KPConv [7] block to extract point features in an end-to-end manner. Nevertheless, this approach requires substantial computational resources to search neighborhoods in 3D space, which hinders them from real-time processing. In this work, we partition points into frustum regions and extract local features of points in a 2D manner, avoiding time-consuming neighbor searches. Based on the learned point features, FRNet predicts results over the entire point cloud in an end-to-end manner with high efficiency.
@@ III-A Pilot Study
LiDAR segmentation aims to assign semantic labels to each point in the point cloud [62]. A common approach is to operate directly on the LiDAR points without any pre-processing, following the paradigm established by PointNet++ [63]. However, point-based operations often require extensive computations and can yield sub-optimal performance due to the sparsity of outdoor scenes. To address this problem, many recent works have proposed projecting LiDAR points onto a regular intermediate representation, known as range images, based on spherical coordinates to enhance efficiency. Specifically, given a point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}$ with $N$ points, where $\mathbf{p}_{i}\in\mathbb{R}^{3+L}$ includes the coordinate $(x_{i},y_{i},z_{i})$ and additional $L$-dimensional features (intensity, elongation, etc.), we map the point to its corresponding 2D position $(u_{i},v_{i})$ in the range image via the following transformation:
$\binom{u_{i}}{v_{i}}=\binom{\frac{1}{2}[1-\arctan(y_{i},x_{i})\pi^{-1}]W}{[1-(\arcsin(z_{i}d_{i}^{-1})+\phi_{\text{down}})\phi^{-1}]H}~{},$ (1)
where $d_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}+z_{i}^{2}}$ is the depth of the point; $\phi=|\phi_{\text{up}}|+|\phi_{\text{down}}|$ represents the vertical field-of-view (FOV) of the sensor, with $\phi_{\text{up}}$ and $\phi_{\text{down}}$ being the inclination angles in the upward and downward directions, respectively; $H$ and $W$ are the height and width of the range image. However, such projection often results in the loss of occluded points and requires post-processing techniques to reconstruct the entire semantic information [36].
We conducted a pilot study on various range-view methods, including RangeNet++ [34], SalsaNext [64], FIDNet [59], and CENet [61]. We compared their performances on the original points (i.e., after the 2D-to-3D projection) using the widely-used KNN post-processing with different nearest neighbor parameters, as well as their performance on the 2D range images. As illustrated in Fig. 2, the post-processing procedures are often unsupervised, and the choice of hyperparameters can significantly impact the final performance. To address this issue, we propose a scalable Frustum-Range Network (FRNet) capable of directly predicting semantic labels for each point while maintaining high efficiency.
@@ III-B Frustum-Range Architecture
Problem Definition. Given a point cloud $\mathcal{P}$ acquired by the LiDAR sensor, the objective of FRNet is to employ a feed-forward network $\mathcal{G}$ to predict semantic labels $\mathcal{\widehat{Y}}$ for each individual point as follows:
$\mathcal{\widehat{Y}}=\mathcal{G}(\mathcal{P},\theta)~{},$ (2)
@@
Frustum Feature Encoder. Accurate prediction of semantic labels for each point requires the extraction of individual features. Before converting the point cloud into a 2D frustum representation, the frustum feature encoder plays a pivotal role in extracting per-point features through a series of MLPs, which is crucial for accurate label prediction. The frustum feature encoder follows three steps. ① We divide the point cloud into $H$ and $W$ frustum regions along the vertical and horizontal directions, respectively. According to Eq. 1, for each point $\mathbf{p}_{i}$ in the point cloud $\mathcal{P}$, we calculate its frustum position $(u_{i},v_{i})$. Points sharing the same $(u,v)$ coordinates are grouped into the same frustum region. This results in a set of $M$ frustum regions, defined as $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{M}\}$, where $M=H\times W$ and $\mathcal{P}_{i}\in\mathbb{R}^{N_{i}\times(3+L)}$ consists of $N_{i}$ points within the $i$-th frustum region. ② To explicitly embed the frustum structure, we leverage an average pooling function within each frustum region to obtain the cluster points $\mathcal{\widetilde{P}}=\{\mathcal{\widetilde{P}}_{1},\mathcal{\widetilde{P}}_{2},...,\mathcal{\widetilde{P}}_{M}\}$, where $\mathcal{\widetilde{P}}_{i}\in\mathbb{R}^{1\times(3+L)}$. Subsequently, we construct a frustum graph between the point cloud $\mathcal{P}$ and the cluster points $\mathcal{\widetilde{P}}$ within each frustum region. The per-point features are embedded as follows:
@@
[…]
To efficiently extract context information for the points, we employ a frustum-to-point fusion block to integrate local features from the frustum region into the points. First, ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is back-projected to the point-level features according to the projection index. Subsequently, the back-projected features are concatenated with $\mathcal{F}_{p}^{i-1}$ and passed through an MLP to update the per-point features, avoiding high computational costs from neighborhood searches:
[…]
where $\text{Conv}(\cdot)$ represents a convolution layer with batch normalization and an activation function. Finally, the fused features are utilized to further enhance the frustum features in a residual-attentive manner:
[…]
@@
(a) Scene 1
(b) Scene 2
(c) Mixed Scene
Figure 5: FrustumMix illustration. (a) and (b) show the original two LiDAR scenes. (c) presents the mixed scenes generated by the FrustumMix strategy, where scene 1 is colored green and scene 2 is colored purple.
(a) Before RangeInterpolation
(b) After RangeInterpolation
Figure 6: RangeInterpolation illustration. (a) shows the raw range image with numerous empty pixels. (b) illustrates the range image after applying RangeInterpolation, resulting in a smoother and more compact representation.
Fusion Head Module. FRNet aims to assign a category to each point in the point cloud, which requires point-level features for end-to-end label prediction. Although the output point features from the last layer of the frustum-point fusion module can yield promising performance, we observe this is not the optimal solution. Features from different levels have different context-aware information. Lower-level features are geometric while higher-level features are semantically rich. Leveraging these different level features can enhance performance. Drawing inspiration from FIDNet [59] and CENet [61], we first combine all point and frustum features from each stage of the backbone and fuse them to reduce channel dimension:
@@
where $V$ is the number of stages in the backbone. Fusing backbone features with different receptive fields helps obtain more context-aware and semantic information. Additionally, point features $\mathcal{F}_{p}^{\text{out}}$ contains local information from neighborhood frustum regions, while frustum features $\mathcal{F}_{f}^{\text{out}}$ contains global information within the corresponding frustum region. Meanwhile, $\mathcal{F}_{p}^{0}$ from the frustum feature encoder module constructs graph structures within the frustum region and between neighborhood frustum regions. Progressively fusing these features helps extract features from local to global, and from geometric to semantic, resulting in more accurate predictions. To fuse $\mathcal{F}_{f}^{\text{out}}$ and $\mathcal{F}_{p}^{\text{out}}$, we first back-project the frustum features to point-level and use an MLP layer to ensure the same channels as $\mathcal{F}_{p}^{\text{out}}$. Then, an addition operation is applied to fuse the two features. The same operation is utilized to fuse the new features with $\mathcal{F}_{p}^{0}$. The overall operation can be formulated as:
@@
Optimization. In the conventional setup, a point-level loss function is commonly employed to supervise the final semantic scores. However, in this work, we propose optimizing frustum features with additional auxiliary heads, utilizing the frustum-level loss function. This approach requires generating pseudo frustum labels. Typically, a frustum region encompasses multiple points, which may not all belong to the same category. To address this, we adopt a statistical approach inspired by sparse-voxel methods [16]. We generate a pseudo label for each frustum by first counting the frequency of each category existing in the given frustum region, and then choosing the category with the highest frequency as the pseudo label for the frustum. Finally, the frustum-level features are supervised using the generated pseudo labels. The overall loss function can be formulated as follows:
@@
where $\mathcal{L}_{p}$ and $\mathcal{L}_{f}$ represent the point-level and frustum-level loss functions, respectively, and $\lambda$ is a hyperparameter that controls the weight of frustum supervision. $\mathcal{L}_{p}$ is calculated by cross-entropy loss, while $\mathcal{L}_{f}$ consists of cross-entropy loss, Lovász-Softmax loss [65], and boundary loss [66].